a gas has a volume of 4.2 l at a temperature of 243. K and pressure of 803 mmHg. what is the new temperature (in celsius) if the volume becomes 10.55 L when the pressure drops to 0.37 atm


1
Expert's answer
2021-04-29T07:18:18-0400

P1V1/T1 = P2V2/T2

803 × 0.0013158

= 1.057 atm

= 1.057 × 4.2/243 = 0.37 × 10.55/T2

0.0183 = 3.9035/T2

0.0183 T2 = 3.9035

T2 = 213.31k

213.32-273 = -59.69°C